x Company prepares annual financial statements. On September 1, 2017, X Company paid $66,000 in advance for a ear insurance policy. After the adjusting entry on December 31, 2017, what will X Companys 2017 financial statements Prepaid Insurance, $16,50; Insurance expense, $49,500 Prepaid Insurance, $49,500; Insurance expense, $16,500 Prepaid Insurance, $55,000; Insurance expense, $11,000 Prepaid Insurance, $1,000; Insurance expense, $55,000 Prepaid Insurance, $66,000; Insurance expense Prepaid Insurance, S0; Insurance expense, 866,000 , So
Answer #1

Prepaid Insurance on september 1, 2017 = $66,000

Insurance expired each month = 66000 / 24 = $2,750

Insurance expired in 2017 = 2750 * 4 = $11,000.

X Company will report in finacial statements for 2017:

c. Prepaid Insurance $55,000; Insurance expense $11,000.